# OpenVision 3  [Uncaptioned image]: A Family of Unified Visual Encoder for Both Understanding and Generation

### v1

@@ 4.2 Reconstruction performance
Figure 3: Loss visualization with only reconstruction loss. We trained our tokenizer with and without the understanding loss, respectively. In Figure (a), the inclusion of semantic loss leads to a lower image reconstruction loss, suggesting that semantic supervision can, in turn, enhance reconstruction performance. Figures (c) and (d) reveal that both caption and contrastive losses decrease even without explicit semantic training, further demonstrating that the two objectives are mutually beneficial.
[…]
SD-VAE DiT 2.27 278.2 0.83 0.57
SD-VAE SiT 2.06 270.3 0.82 0.59
UniTok LlamaGen 2.51 216.7 0.82 0.57
CLIP RAE 2.54 256.4 0.80 0.54
OpenVision RAE 2.44 262.2 0.80 0.53
OpenVision 3 RAE 1.89 289.2 0.84 0.59
Table 4: Class-conditional image generation on ImageNet 256x256. We report gFID, Inception Score (IS), Precision (Pre.), and Recall (Rec.).
4.3 Generation performance
As shown in we report generation Fréchet inception distance (gFID), Inception Score (IS), Precision (Pre.), and Recall (Rec.) as evaluation metric. We present the generative performance of each tokenizer when paired with its respective compatible generator. For low-level tokenizer, we evaluate SD-VAE with traditional diffusion-based generative models (DiT and SiT) (Peebles and Xie, 2023; Ma et al., 2024). For semantic tokenizer, we select CLIP and RAE generator for fair comparison with our tokenizer.According to Table 4OpenVision 3 outperforms these tokenizers across all the metrics. For example, we achieve better gFID when compared to SD-VAE with improved generator SiT(1.89 vs. 2.06). Our tokenizer also surpasses semantic encoders like CLIP (Radford et al., 2021) by a large margin in generation(1.89 vs. 2.54).
Visualization.
As shown in Figure 4, we visualize the generated results with RAE and OpenVision 3, which show the strong capability of our tokenizer in generating high-quality samples with great fidelity and diversity.
Figure 4: Qualitative results of class-conditional ImageNet-256 generation. Under the RAE framework, our OpenVision 3 is able to generate high quality images.
4.4 Understanding performance
To evaluate the semantic representation capability of OpenVision 3, we integrate it into the LLaVA-1.5 framework and conduct training following its standard training configurations. Due to the fixed downsample size of VAE, we keep the same encoded token numbers with OpenAI CLIP for fair comparison. In Table 3, we compare our tokenizer with CLIP and present the results on multiple multimodal benchmarks, including MME (Fu et al., 2023), ScienceQA (Saikh et al., 2022), SeedBench (Li et al., 2024a), GQA (Hudson and Manning, 2019) and POPE (Li et al., 2023c). According to the table, OpenVision 3 can match or exceed the understanding performance of CLIP on general multimodal tasks. For example, our tokenizer consistantly surpasses CLIP on SeedBench (62.4 vs. 62.2 and 66.0 vs. 65.4) and POPE (83.7 vs. 82.9 and 85.3 vs. 84.7). It can be observed that our unified tokenizer is comparable to the understanding-oriented CLIP in terms of semantic comprehension, and even demonstrates a clear advantage in certain aspects.
4.5 Interaction of understanding and reconstruction
@@
In Figure 2, we remove reconstruction loss and train only with semantic loss. The blue curve represents the baseline loss, while the red curve denotes the model trained without the reconstruction loss. According to the loss curves in Figure 2(a) and Figure 2(b), even in the absence of reconstruction objectives, the reconstruction loss still exhibits a substantial decline, suggesting that our semantic objectives contribute significantly to image reconstruction. Furthermore, comparing the red and blue curves in Figure 2(c) and Figure 2(d), it is evident that the incorporation of the reconstruction loss leads to no significant change in either caption or contrastive loss. These observations collectively indicate a mutually beneficial synergy between the two types of losses.
@@
In Figure 3, we remove understanding loss and train only with reconstruction-driven signals. The red curves here denote the loss without reconstruction-driven signals.  Figure 3(c) and Figure 3(d) show that in the absence of semantic supervision, the contrastive loss remains almost stagnant, whereas the caption loss exhibits a marginal decline. This indicates that the reconstruction task intrinsically facilitates semantic tasks that are also generative in nature. Moreover, as seen in Figure 3(a), the addition of semantic loss paradoxically improves reconstruction performance, providing further evidence of the synergistic relationship between these two branches.
@@ 5 Conclusion
This work introduces OpenVision 3, a unified vision encoder for both understanding and generation. We innovatively couple a VAE with a ViT to form a unified architecture, and generate a single, unified representation for different downstream tasks. For the efficient training of our tokenizer, we propose a new training paradigm with both reconstruction- and semantics-driven signals for joint learning. Comprehensive evaluations reveal that our model yields superior results across generative and understanding tasks through low-cost training. OpenVision 3 outperforms current other unified tokenizer in reconstruction and generation, and shows competitive ability with CLIP on semantic tasks. To facilitate future research in the community, we will fully open-source our training code, data, and tokenizer checkpoints.
Acknowledgements
We would like to thank TPU Research Cloud (TRC) program and Google Cloud Research Credits program for supporting our computing needs.

### v2

@@ 4.4 Understanding performance
Figure 5: Loss visualization with only reconstruction loss. We trained our tokenizer with and without the understanding loss, respectively. In Figure (a), the inclusion of semantic loss leads to a lower image reconstruction loss, suggesting that semantic supervision can, in turn, enhance reconstruction performance. Figures (c) and (d) reveal that the caption loss exhibits a slight downward trend despite the lack of direct semantic signals, and the contrastive loss remains almost stagnant.
5 Discussion
5.1 Reciprocal synergy between understanding and reconstruction
@@
In Figure˜4, we remove reconstruction loss and train only with semantic loss. The blue curve represents the baseline loss, while the red curve denotes the model trained without the reconstruction loss. According to the loss curves in Figure˜4(a) and Figure˜4(b), even in the absence of reconstruction objectives, the reconstruction loss still exhibits a substantial decline, suggesting that our semantic objectives contribute significantly to image reconstruction. Furthermore, comparing the red and blue curves in Figure˜4(c) and Figure˜4(d), it is evident that the incorporation of the reconstruction loss leads to no significant change in either caption or contrastive loss. These observations collectively indicate a mutually beneficial synergy between the two types of losses.
@@
In Figure˜5, we remove understanding loss and train only with reconstruction-driven signals. The red curves here denote the loss without reconstruction-driven signals.  Figure˜5(c) and Figure˜5(d) show that in the absence of semantic supervision, the contrastive loss remains almost stagnant, whereas the caption loss exhibits a marginal decline. This indicates that the reconstruction task intrinsically facilitates semantic tasks that are also generative in nature. Moreover, as seen in Figure˜5(a), the addition of semantic loss paradoxically improves reconstruction performance, providing further evidence of the synergistic relationship between these two branches.
Table 6: Effectiveness of VAE in reconstruction. We study the influence of VAE latents in reconstruction performance. Compared to the variant without a VAE, OpenVision 3 demonstrates a significant advantage in rFID, achieving superior reconstruction quality. Our tokenizer maintains competitive results on PSNR, SSIM, and LPIPS with slight differences.
[…]
w/o VAE 32.82 0.935 0.060 0.980
OpenVision 3 30.33 0.885 0.061 0.216
Table 7: Effectiveness of VAE in Generation. Under the VAE latent space, OpenVision 3 outperforms its counterpart by 1.23 gFID, showing better generation ability.
[…]
w/o VAE 9.68
OpenVision 3 8.45
Table 8: Understanding performance comparison of unified tokenizer with and without VAE. We evaluate the multimodal understanding capabilities of OpenVision 3 and its non-VAE counterpart under the LLaVA-1.5 and LLaVA-NeXT architectures. Under identical training settings, it can be observed that OpenVision 3 consistently outperforms the non-VAE version. Under the LLaVA-1.5 framework, our tokenizer outperforms the counterpart in four out of six metrics. Notably, under the LLaVA-NeXT framework, our model achieves a comprehensive lead across all metrics.
LLaVA-1.5 Framework
OpenVision 3 VAE + B/2 196 224 1382 287 62.4 73.0 58.0 83.7
w/o VAE B/16 196 224 1366 307 62.2 72.7 57.6 83.8
LLaVA-NeXT Framework
OpenVision 3 VAE + B/2 196 224 1383 348 63.6 70.5 59.1 84.4
w/o VAE B/16 196 224 1346 296 62.0 67.8 57.7 83.8
((a))
((b))
((c))
((d))
Figure 6: Loss visualization with different ViT decoder sizes. We use M/1, B/1 and L/1 variants as ViT decoder and draw the loss curves. As shown in Figures (a) and (b), employing a Large-sized decoder leads to significant training instability. While varying the decoder size has negligible impact on understanding performance, the Base-sized decoder effectively reduces reconstruction loss while maintaining training stability.
5.2 VAE latents help unified learning
A core design of our tokenizer involves feeding VAE latents into a ViT for encoding. VAEs are originally designed for low-level visual encoding to capture pixel-level image features. However, we found that performing unified modeling with VAE latents substantially enhances image generation capabilities without compromising multimodal understanding performance, as shown in $Tab.˜7$ and Tab.˜8. We argue that high-quality VAE latents enable a single ViT encoder to generate a single, unified visual representation that supports both generation and understanding. This avoids the need for separate encoding processes followed by concatenation or feature fusion. Furthermore, within this VAE latent space, we can achieve superior alignment between features for understanding and generation, leading to high performance in both capabilities.
To evaluate the necessity of VAE, we conducted a comparative analysis by training 2 tokenizers with raw image tokens and VAE latents, respectively. Specifically, since the downsample rate of FLUX-VAE is $8$, we separately train VAE+ViT-B/2 and ViT-B/16 as the visual tokenizers to keep the same image token number. Apart from the difference in the encoder architecture, all other structural components and hyperparameters were kept strictly identical. For both sides, we pretrain at $\times128$ resolution for 1000 epochs, and then finetune for 200 epochs at $\times256$ for generation, and $\times224$ for understanding.
In Tab.˜7, we report the reconstruction performance with and without VAE. Compared to the standalone ViT, our tokenizer outperforms it by a large margin in rFID (0.216 vs. 0.980), demonstrating our high-quality reconstruction results. Meanwhile, there are marginal gaps in the other three metrics: PSNR, SSIM, and LPIPS. Furthermore, we train both tokenizers under RAE framework with 100 epochs to study the effectiveness of VAE in generation. As shown in Tab.˜7, our model achieves a gFID of 8.45 on ImageNet with the VAE, representing a 1.23 improvement over the 9.68 gFID obtained without it. In Tab.˜8, we evaluated the understanding capabilities of the tokenizer when the VAE is removed under both LLaVA-1.5 and LLaVA-NeXT frameworks. The results show a noticeable performance gap compared to OpenVision 3. Our tokenizer outperforms the baseline in 4 out of 6 metrics under the LLaVA-1.5 setting; notably, when using the stronger LLaVA-NeXT baseline, our advantage becomes even more pronounced. This suggests that although the VAE primarily extracts low-level visual features, it also contributes to the enhancement of multimodal understanding capabilities. These findings demonstrate that unifying tasks in the VAE latent space yields simultaneous gains in reconstruction fidelity, generative quality, and comprehensive understanding, validating the necessity of the VAE in our unified learning framework.
5.3 The selection of ViT decoder size
In the reconstruction branch, we use a ViT and a linear layer as the pixel decoder. In Figure˜6, we investigate the impact of the size of the ViT decoder on reconstruction and understanding. Varying the ViT decoder variants for pixel-level decoding has minimal impact on the understanding branch. As illustrated in Figure˜6(c) and Figure˜6(d), the semantic signals remains nearly invariant across different ViT decoder sizes. Regarding reconstruction learning, increasing the decoder size from M/1 to B/1 results in a significant reduction in pixel-level loss, as shown in Figure˜6(a). However, employing the larger L/1 variant leads to highly unstable training, as depicted by the brown curve. In terms of training time, we conduct experiments for 200 epochs for each size. The total training durations for the M/1, B/1, and L/1 decoder on a TPU v4-256 pod are 6.1, 6.0, and 6.9 hours, respectively. Taking into account both performance and computational efficiency, employing Base-sized ViT as the decoder strikes an optimal balance.
5.4 The selection of ViT encoder size
In Tab.˜9, we explore the scaling behavior of the ViT encoder on the image reconstruction and generation. Generation results are evaluated under RAE pipeline with 100 training epochs. The results indicate that B-size and L-size tokenizers yield comparable results in reconstruction and generation. In contrast,  Tab.˜4 and Tab.˜5 demonstrate that scaling up to L-size provides substantial gains in multimodal understanding. We argue that although increasing the size leads to a more representative tokenizer naturally, the outer VAE architecture imposes a theoretical ceiling on the generative performance. This bottleneck marginalizes the gains derived from larger encoders, presenting an open problem that requires further investigation in future research.
Table 9: Effectiveness of encoder size in reconstruction and generation. B size and L size show similar performance in reconstruction and generation ability.
Model Reconstruction Generation
[…]
OpenVision 3-B 30.92 0.902 0.053 0.187 8.45 144.7 0.78 0.36
OpenVision 3-L 30.96 0.902 0.052 0.186 8.89 148.3 0.78 0.36
@@ 6 Conclusion
This work introduces OpenVision 3, a unified vision encoder for both understanding and generation. We innovatively couple a VAE with a ViT to form a unified architecture, and generate a single, unified representation for different downstream tasks. For the efficient training of our tokenizer, we propose a new training paradigm with both reconstruction- and semantics-driven signals for joint learning. Comprehensive evaluations reveal that our model yields superior results across generative and understanding tasks. OpenVision 3 outperforms current other unified tokenizer in reconstruction and generation, and shows competitive ability with CLIP on semantic tasks. Furthermore, our architecture enables a mutual promotion relationship between image understanding and generation. To facilitate future research in the community, we will fully open-source our training code, data, and tokenizer checkpoints.